Expand each logarithmic expression as much as possible:
log sub b (x^3y/z^2)
I've been trying to work at this and another all day but can't figure it out to save my algebra butt. final is tomorrow and she said to expect ones like this. Any help would be just so wonderful.

Answer by jcmtnez(53)


log (x^3y/z^2) = 3log(x)+log(y)-2log(z)
Explanation
The logarithm of the product of two number is equal to the sum of the logarithm of each number.
log(xy)=log(x)+log(y)
The logarithm of the division of two numbers is equal to the substraction of the logarithm of each number.
log(x/y)=log(x)-log(y)
Remember that the base does not change. You may use this properties to solve any problem like the above.